1) Calculate q, w, Д ed--if 2.25 mol of an ideal gas with Cm 3R/2 undergoes a reversible adiabatic expansion from an initial volume Vi - 5.50 m* to a final vol- ume Vf 25.0 m3. The initial temperature is 275 K.
Answer #1

TV^(gamma-1)=constant

Gamma=Cp/Cv=5/3

=>275×(5.5)^([5/3]-1) =T2×(25)^[(5/3)-1]

=>T2= 100.2K-4904